### v1

@@ 4 Triangle Counting via Mobile Agents
In our algorithm, the agents (technically representing the nodes they are sitting at) first scan their neighbourhood. Once all the information about the neighbourhood is collected, the agents now count the number of common neighbourhoods between two adjacent agents. After each agent receives that count, the sum of each such count stored at each agent is evaluated, which when divided by three gives us the number of triangles in $G$. The algorithm runs in three phases. In the first phase, the agents learn their neighbours. In the second phase, the agents check the number of common neighbours with each of its adjacent agents. In this phase, each agent $r_{i}$ also counts the number of local triangles with $r_{i}$ as a vertex and the number of triangles with $(r_{i},r_{j})$ as an edge, where $r_{j}$ is an adjacent agent to $r_{i}$. In the third and last phase, each agent collects the local triangle count from every other agent and counts the number of triangles in $G$. We now explain each phase in detail, below.
4.1 Phase 1: Know the Neighbourhood
Each node is occupied by a distinct (via their IDs) agent before the start of the algorithm and we represent each node of $G$ by its stationed agent. The algorithm starts with Phase 1, where each agent discovers and records its neighbour. As the IDs of the agents are unknown, agents cannot synchronously start scanning the neighbourhood. With each agent executing the same algorithm, agents may not find other agents at their exact place when they move. Therefore, we need to ensure that each gets to record the correct neighbour set of it. To do this we exploit the ID bits of the agents. We use the fact the IDs of the agents are distinct. To this end, we state the following lemma.
Lemma 1.
Let $r_{i}$ and $r_{j}$ be two distinct agents in $\mathcal{R}$, with their IDs field being $r_{i}.ID$ and $r_{j}.ID$ respectively. Then, there exists at least one dissimilar bit in $r_{i}.ID$ and $r_{j}.ID$ with one being 00 and the other being $1$.
@@ 4.1 Phase 1: Know the Neighbourhood
Let $r_{j}$ (placed at node $v$) be a neighbour of $r_{i}$ (at node $u$). Then there exists a port joining $u$ to $v$ having port number between $[0,\delta(u)-1]$. Since $r_{i}$ and $r_{j}$ have distinct IDs, there exists a bit (say the $p^{th}$ bit from the right) in $r_{i}.ID$ and $r_{j}.ID$, which are different from one another (one bit being 00, the other being $1$)[Lemma 1]. Without the loss of generality, let us assume that the $p^{th}$ bit of $r_{i}$ is $1$ and the $p^{th}$ bit of $r_{i}$ is 00. Therefore, when the $p^{th}$ bit becomes the current ID bit, $r_{i}$ starts exploring its neighbours one by one during which it finds the agent $r_{j}$ now stationary at $v$ with current ID bit 00. Therefore $r_{i}$ records $r_{j}$ as its neighbour (simultaneously $r_{j}$ also registers $r_{i}$ in its neighbour list). In a similar way, $r_{i}$ records all of its neighbours exhaustively either by visiting a stationary agent in its neighbourhood or by meeting an exploratory agent from its neighbour. Phase 1 runs for $\log\lambda$ rounds with each round consisting of $O(\Delta)$ sub-rounds to allow the agents to visit each neighbour. Therefore, Phase 1 completes in $O(\Delta\cdot\log\lambda)$ rounds. ∎
Therefore, with the end of Phase 1, each agent has now enlisted its neighbouring agents in its memory. The algorithm now moves to Phase 2 where each agent $r_{i}$ counts the number of triangles with $r_{i}$ as one of its vertex (node).
@@ 4.2 Phase 2: Local Triangle Counting
In phase two, each agent, now equipped with the list of its neighbouring agents, visits each of its neighbours once again, in exactly the similar fashion described in Phase 1. Whenever the agent $r_{i}$ meets its neighbour $r_{j}$, they communicate to find out the number of common neighbours they both have. Here as the nodes are anonymous, they are identified using the mobile agents that reside on the nodes. This communication is also used the update the following variables of $r_{i}$:
@@
    $r_{i}.edge(r_{j}):$ A variable which stores the count of the number of common neighbours of $r_{i}$ and $r_{j}$ which represents the number of triangles based on the edge containing $(r_{i},r_{j})$. $r_{i}$ has $\Delta$ such variables $r_{i}.edge(x)$, where $x$ represents another agent that has an edge with $r_{i}$. Each of these variables are initially set to 00. A variable $r_{i}.edge(x)$ remains 00, if there are no triangles with $(r_{i},x)$ as an edge.
@@
    $r_{i}.local\_sum:$ Initially set to 00, adds up the counts of the number of common neighbours for each distinct neighbour of $r_{i}$. As $r_{i}$ finds its neighbours one by one, it cumulatively adds up the count of the number of common neighbours with each of its neighbours. Mathematically, $r_{i}.local\_sum=\Sigma_{j}r_{i}.edge(r_{j})$, where the sum runs over every neighbouring agent $r_{j}$ of $r_{i}$.
@@
First, we see that if the agents $r_{i}$ and $r_{j}$ have a common neighbour $r_{k}$, then, $(r_{i},r_{j},r_{k})$ form a triangle with $r_{i}$ as one of the vertex. Now, the triangle $(r_{i},r_{j},r_{k})$ has been tallied once while counting the common neighbour of $r_{i}$ with $r_{j}$ and re-counted again as a common neighbour of $r_{i}$ with $r_{k}$. Therefore, each $(r_{i},r_{j},r_{k})$ with $r_{i}$ as vertex, is counted twice, once through $r_{j}$ and other time through $r_{k}$. So, the number of triangles with agent $r_{i}$ as a vertex is exactly half of $r_{i}.local\_sum$. ∎
Therefore at the end of Phase 2, each agent $r_{i}$ gets a local count of the number of triangles with $r_{i}$ as a vertex and a list of the number of triangles that are based on the edges adjacent to $r_{i}$. In the next phase, the algorithm finds the number of triangles of $G$ using the counts generated in this phase.
@@ 4.3 Phase 3: Counting the Number of Triangles in $G$
To find the number of triangles in $G$, we need to take into account the local triangle counts of each agent. We first need to accumulate the $r_{i}.local\_sum$ values of each of the $n$ agents and calculate the triangles of $G$ from there. However, due to the absence of a $leader$ agent and the difficulty of synchronising the movement of the agents, gathering the $r_{i}.local\_sum$ is not straightforward. The high-level idea is to communicate repeatedly with the neighbours and continually gather up the values of $local\_sum$ of each agent hop by hop till each agent gets every $local\_sum$ value.
As described in the previous case, the agents can assuredly communicate with all their neighbours in $\Delta\cdot\log\lambda$ rounds. Inside the first $\Delta\cdot\log\lambda$ round of this phase, the agents communicate with their neighbours and exchange the variable $local\_sum$. An agent $r_{i}$, along with its own $r_{i}.local\_sum$, collects $r_{j}.local\_sum$ values for each of its neighbour $r_{j}$ and stores it in the memory. Within the next $\Delta\cdot\log\lambda$ rounds, $r_{i}$ meets its neighbour agents $r_{j}$ again to check if they have collected any new $local\_sum$ values (possibly from their own neighbour). $r_{i}$ again stores any new $local\_sum$ record (of its neighbour’s neighbour) that it receives through its neighbour $r_{j}$ in the current phase. By the end of the second phase, $r_{i}$ has the information of the $local\_sum$ of all the agents that are at a distance of $2-hops$ from it. Now, with the completion of $D$ (the diameter of the graph $G$) such $\Delta\cdot\log\lambda$ rounds, each agent $r_{i}$ now has the $local\_sum$ record of every agent in $G$. The number of triangles in $G$ is finally calculated by each agent by summing up the $n$ $local\_sum$ values that they have gathered and dividing it by $6$.
Lemma 4.
@@
Since $r_{i}$ and $r_{k}$ are at a distance of $d$ hops from each other, there exists a sequence of agents (nodes) $(r_{i},r_{i_{1}},r_{i_{2}},\dots,r_{i_{d-1}},r_{k})$ from $r_{i}$ to $r_{k}$. In the first $O(\Delta\cdot\log\lambda)$ rounds, $r_{i}$ can communicate with $r_{i_{1}}$ to get the value of $r_{i_{1}}.local\_sum$. In the meanwhile, $r_{i_{1}}$ also collects the value of $r_{i_{2}}.local\_sum$ in the same $O(\Delta\cdot\log\lambda)$ round. Therefore, in the second sub-phase of $O(\Delta\cdot\log\lambda)$ rounds, when $r_{i}$ communicates with $r_{i_{1}}$ again, $r_{i}$ receives the value of $r_{i_{2}}.local\_sum$ through $r_{i_{1}}$. Continuing in a similar way, $r_{i}$ is guaranteed to receive the value of $r_{k}.local\_sum$ from $r_{k}$ at a distance of $d$ hops through the agents $r_{i_{1}},r_{i_{2}},\dots,r_{i_{d-1}}$ by the end of $d\cdot O(\Delta\cdot\log\lambda)$ rounds. ∎
Let $D$ denote the diameter of the graph $G$. Since any two agents in $G$ are located at a distance of at most $D$ hops, the following lemma follows from Lemma 4.
Lemma 5.
After the end of Phase 3, which takes $O(D\cdot\Delta\cdot\log\lambda)$ rounds, each agent $r_{i}$ has the complete record of $local\_sum$ values of every agent in $G$. Here, $D$ is the diameter of the graph $G$.
If the diameter $D$ of the graph is unknown, $D$ could be replaced by $n$.
Lemma 6.
@@
Notation: We shall denote the number of triangle containing the vertex(node) $v$ and the total number of triangles in $G$ with $\mathbf{T}(v)$ and $\mathbf{T}(G)$, respectively. For a node $v$ with a robot $r_{i}$ on it, we use $r_{i}$ and $v$ interchangeably to denote the node.
At the end of this phase, each agent has the value of the number of triangles in $G$ i.e., $\mathbf{T}(G)$. In the following theorem, we assemble the list of results we got during the 3 phases.
@@
    Each agent $r_{i}$ can calculate the number of triangles in $G$, $\mathbf{T}(G)$, in $O(D\cdot\Delta\cdot\log\lambda)$ rounds.
@@ 5 Applications: Truss Decomposition, Triangle Centrality and Local Clustering Coefficient
In this section, we show some novel applications of our triangle-counting algorithm via mobile agents in an anonymous graph.
@@ 5.1 Truss Decomposition
In our algorithm, we use the popular truss decomposition method to first calculate, what is called, $trussness$ value for each edge. The $trussness$ values are then used to construct the $k$-truss sub-graph for any value of $k$.
@@
Algorithm for Truss Decomposition: Let us now propose an algorithm for mobile agents to evaluate $trussness$ for each edge of the graph $G(V,E)$. The algorithm is based on the parallel truss decomposition described in [55]. The $trussness$ values determine a partition (thus an equivalence relation) on $E$, where each class have the edges of $G$ with the same $trussness$ value. Let $\mathbf{t_{k}}$ denote the equivalence class of edges having $trussness=t_{k}$. To find the $k$-truss, $T_{k}$ we construct a sub-graph of $G$ with edges from the equivalence classes $\mathbf{t_{k}},\mathbf{t_{k+1}},\dots,\mathbf{t_{max}}$, where $t_{max}$ is the maximum $trussness$ of any edge in $G$. Mathematically, $T_{k}=\cup_{i=k}^{t_{max}}\mathbf{t_{i}}$. Therefore by computing the trussness for each edge in $G(V,E)$, we obtain a partition (equivalence classes) of $E$, thereby obtaining the $k$-trusses of $G$ for any $k$ by taking the union of the equivalent classes. Therefore, the $k$-truss decomposition of a graph is equivalent to computing the $trussness$ of each edge in the graph.
$trussness$ values of each edge in $G$ can be computed efficiently using existing serial and parallel algorithms. The serial truss decomposition algorithm first calculates $support$ for each edge and iteratively removes a single edge each time until all the edges of the graph are removed. Each time, the edges are sorted in ascending order and the edge with the lowest $support$ is removed. The removed edge (say $e$) keeps its final $support+2$ as its $trussness$ value. Once the $e$ is removed, the $support$ value of the edges that formed the triangle with $e$ is re-evaluated. The remaining edges are once again sorted in order before running another pass. The need to sort the edges in order of $support$ makes the algorithm inherently sequential.
In the parallel version of the algorithm, the sorting condition is relaxed with the use of $h-index$ updating. For a set of integers $S$, the $h-index$ of $S$ is defined as the largest number $h$ such that there are at least $h$ elements in $S$ that are equal to and greater than $h$. The $trussness$ of an edge is related to $h-index$, for, the $trussness$ of an edge $e$ can be thought of as the largest $k$ such that it is contained in at least $k$ triangles whose edges have a $trussness$ value of at least $k$. In the algorithm, each edge $e$ in $G$ is initialized to its $support$ as the first approximation to its $trussness$. Now, the $support$ values of all triangles with $e$ as an edge are stored in a set $L$ and its $h-index$ is computed. At each iteration, the $h-index$ of $e$ is updated to the smallest of its current value and the $h-index$ of $L$. The algorithm iteratively updates an edge’s $h-index$ by computing the $h-index$ of all edges that support it, until achieving convergence when no updates would happen. The final $h-index$ of each edge before, no further updates happen, provide the $trusness$ value for each edge.
@@
    $r_{i}.edge\_set$ to store the edges incident on a particular node containing $r_{i}$. $r_{i}$ stores only edges $(r_{i},r_{j})$, where $r_{j}>r_{i}$. $edge\_set$ requires $O(\Delta)$ bit of memory, where $\Delta$ is the highest degree of a node in $G$. The ordering of edges ensures that each edge becomes associated with exactly one agent.
@@
    $L$ also associated with each edge in $edge\_set$. $L$ is used to store the values $\min\{h(e^{\prime}),h(e^{\prime\prime})\}$ for for each triangle $(e,e^{\prime},e^{\prime\prime})$ containing $e$.
@@
    $scheduled$ to identify the edges in the $edge\_set$ that need further updating in consequent rounds of the algorithm. Initially, every edge in $edge\_set$ is scheduled ($true$).

### v2

@@ 4 Triangle Counting via Mobile Agents
In our algorithm, the agents (technically representing the nodes they are sitting at) first scan their neighbourhood. Once all the information about the neighbourhood is collected, the agents now count the number of common neighbourhoods between two adjacent agents. After each agent receives that count, the sum of each such count stored at each agent is evaluated, which when divided by three gives us the number of triangles in $G$. The algorithm runs in three phases. In the first phase, the agents learn their neighbours. In the second phase, the agents check the number of common neighbours with each of their adjacent agents. In this phase, each agent $r_{i}$ also counts the number of local triangles with $r_{i}$ as a vertex and the number of triangles with $(r_{i},r_{j})$ as an edge, where $r_{j}$ is an adjacent agent to $r_{i}$. In the third and last phase, each agent collects the local triangle count from every other agent and counts the number of triangles in $G$. We now explain each phase in detail below.
4.1 Phase 1: Neighbourhood Discovery (Protocol MYN)
Each node is occupied by a distinct (via their IDs) agent before the start of the algorithm, and we represent each node of $G$ by its stationed agent. The algorithm starts with Phase 1, where each agent discovers and records its neighbour. As the IDs of the agents are unknown, agents cannot synchronously start scanning the neighbourhood. With each agent executing the same algorithm, agents may not find other agents at their exact place when they move. Therefore, we need to ensure that each gets to record the correct neighbour set of it. To do this, we exploit the ID bits of the agents. We use the fact the IDs of the agents are distinct. To this end, we state the following lemma.
Lemma 4.1.
Let $r_{i}$ and $r_{j}$ be two distinct agents in $\mathcal{R}$, with their IDs fields being $r_{i}.ID$ and $r_{j}.ID$ respectively. Then, there exists at least one dissimilar bit in $r_{i}.ID$ and $r_{j}.ID$ with one being 00 and the other being $1$.
@@ 4.1 Phase 1: Neighbourhood Discovery (Protocol MYN)
Let $r_{j}$ (placed at node $v$) be a neighbour of $r_{i}$ (at node $u$). Then there exists a port joining $u$ to $v$ having a port number between $[0,\delta(u)-1]$. Since $r_{i}$ and $r_{j}$ have distinct IDs, there exists a bit (say the $p^{th}$ bit from the right) in $r_{i}.ID$ and $r_{j}.ID$, which are different from one another (one bit being 00, the other being $1$)[Lemma 4.1]. Without the loss of generality, let us assume that the $p^{th}$ bit of $r_{i}$ is $1$ and the $p^{th}$ bit of $r_{i}$ is 00. Therefore, when the $p^{th}$ bit becomes the current ID bit, $r_{i}$ starts exploring its neighbours one by one during which it finds the agent $r_{j}$ now stationary at $v$ with current ID bit 00. Therefore $r_{i}$ records $r_{j}$ as its neighbour (simultaneously $r_{j}$ also registers $r_{i}$ in its neighbour list). In a similar way, $r_{i}$ records all of its neighbours exhaustively either by visiting a stationary agent in its neighbourhood or by meeting an exploratory agent from its neighbour. Phase 1 runs for $\log\lambda$ rounds with each round consisting of $O(\Delta)$ sub-rounds to allow the agents to visit each neighbour. Therefore, Phase 1 completes in $O(\Delta\cdot\log\lambda)$ rounds. ∎
Therefore, with the end of Phase 1, each agent has now enlisted its neighbouring agents in its memory. We call this particular step Protocol MYN, a protocol for Meeting-Your-Neighbour. The algorithm now moves to Phase 2 where each agent $r_{i}$ counts the number of triangles with $r_{i}$ as one of its vertex (node).
@@ 4.2 Phase 2: Local Triangle Counting
In Phase 2, each agent, now equipped with the list of its neighbouring agents, visits each of its neighbours once again, in exactly the same manner as described in Phase 1. Whenever the agent $r_{i}$ meets its neighbour $r_{j}$, they communicate to find out the number of common neighbours they both have. Since the nodes are anonymous, they are identified using the mobile agents that reside on the nodes. This communication is also used to update the following variables of $r_{i}$:
@@
    $r_{i}.edge(r_{j}):$ A variable which stores the count of the number of common neighbours of $r_{i}$ and $r_{j}$ which represents the number of triangles based on the edge containing $(r_{i},r_{j})$. $r_{i}$ has $\Delta$ such variables $r_{i}.edge(x)$, where $x$ represents another agent that has an edge with $r_{i}$. This, therefore requires an additional $O(\Delta\cdot\log n)$ memory per agent which forms the dominant factor in the memory requirement. Each of these variables is initially set to 00. A variable $r_{i}.edge(x)$ remains 00, if there are no triangles with $(r_{i},x)$ as an edge.
@@
    $r_{i}.local\_sum:$ Initially set to 00, adds up the counts of the common neighbours for each distinct neighbour of $r_{i}$. As $r_{i}$ finds its neighbours one by one, it cumulatively adds up the count of the number of common neighbours with each of its neighbours. Mathematically, $r_{i}.local\_sum=\Sigma_{j}r_{i}.edge(r_{j})$, where the sum runs over every neighbouring agent $r_{j}$ of $r_{i}$.
@@
First, we observe that if the agents $r_{i}$ and $r_{j}$ have a common neighbour $r_{k}$, then, $(r_{i},r_{j},r_{k})$ form a triangle with $r_{i}$ as one of the vertices. Now, the triangle $(r_{i},r_{j},r_{k})$ has been tallied once while counting the common neighbour of $r_{i}$ with $r_{j}$ and re-counted again as a common neighbour of $r_{i}$ with $r_{k}$. Therefore, each $(r_{i},r_{j},r_{k})$ with $r_{i}$ as vertex, is counted twice, once through $r_{j}$ and the other time through $r_{k}$. So, the number of triangles with agent $r_{i}$ as a vertex is exactly half of $r_{i}.local\_sum$. ∎
Thus, at the end of Phase 2, each agent $r_{i}$ gets a local count of the number of triangles with $r_{i}$ as a vertex and a list of the number of triangles that are based on the edges adjacent to $r_{i}$. In the next phase, the algorithm determines the number of triangles of $G$ using the counts generated in this phase.
@@ 4.3 Phase 3: Counting the Number of Triangles in $G$
To find the number of triangles in $G$, we need to take into account the local triangle counts of each agent. First, we need to accumulate the $r_{i}.local\_sum$ values of each of the $n$ agents and calculate the triangles of $G$ from there. However, due to the absence of a $leader$ agent and the difficulty of synchronising the movement of the agents, gathering the $r_{i}.local\_sum$ is not straightforward. The high-level idea is to communicate repeatedly with the neighbours and continually gather up the values of $local\_sum$ of each agent, hop by hop, until each agent gets every $local\_sum$ value.
As described previously, the agents can reliably communicate with all their neighbours in $\Delta\cdot\log\lambda$ rounds. During the first $\Delta\cdot\log\lambda$ round of this phase, the agents communicate with their neighbours and exchange their $local\_sum$ values. An agent $r_{i}$, along with its own $r_{i}.local\_sum$, collects $r_{j}.local\_sum$ values from each of its neighbours $r_{j}$ and stores them in memory. During the next $\Delta\cdot\log\lambda$ rounds, $r_{i}$ meets its neighbour agents $r_{j}$ again to check if they have collected any new $local\_sum$ values (possibly from their own neighbour). $r_{i}$ again stores any new $local\_sum$ record (of its neighbour’s neighbour) that it receives through its neighbour $r_{j}$ during this phase. By the end of the second phase, $r_{i}$ has the information of the $local\_sum$ of all the agents that are at a distance of $2-hops$ from it. Now, with the completion of $D$ (the diameter of the graph $G$, which is known to the agents) such $\Delta\cdot\log\lambda$ rounds, each agent $r_{i}$ now has the $local\_sum$ record of every agent in $G$. The number of triangles in $G$ is finally calculated by each agent by summing up the $n$ $local\_sum$ values that they have gathered and dividing it by $6$.
Lemma 4.4.
@@
Since $r_{i}$ and $r_{k}$ are at a distance of $d$ hops from each other, there exists a sequence of agents (nodes) $(r_{i},r_{i_{1}},r_{i_{2}},\dots,r_{i_{d-1}},r_{k})$ from $r_{i}$ to $r_{k}$. In the first $O(\Delta\cdot\log\lambda)$ rounds, $r_{i}$ can communicate with $r_{i_{1}}$ to get the value of $r_{i_{1}}.local\_sum$. Meanwhile, $r_{i_{1}}$ also collects the value of $r_{i_{2}}.local\_sum$ during the same $O(\Delta\cdot\log\lambda)$ round. Therefore, in the second sub-phase of $O(\Delta\cdot\log\lambda)$ rounds, when $r_{i}$ communicates with $r_{i_{1}}$ again, $r_{i}$ receives the value of $r_{i_{2}}.local\_sum$ through $r_{i_{1}}$. Continuing in a similar way, $r_{i}$ is guaranteed to receive the value of $r_{k}.local\_sum$ from $r_{k}$ at a distance of $d$ hops through the agents $r_{i_{1}},r_{i_{2}},\dots,r_{i_{d-1}}$ by the end of $d\cdot O(\Delta\cdot\log\lambda)$ rounds. ∎
Let $D$ denote the diameter of the graph $G$. Since any two agents in $G$ are at-most $D$ hops apart, the following lemma follows from Lemma 4.4.
Lemma 4.5.
By the end of Phase 3, which takes $O(D\cdot\Delta\cdot\log\lambda)$ rounds, each agent $r_{i}$ has the complete record of $local\_sum$ values of every agent in $G$.
Lemma 4.6.
@@
Notation: We shall denote the number of triangle containing the vertex (node) $v$ and the total number of triangles in $G$ with $\mathbf{T}(v)$ and $\mathbf{T}(G)$, respectively. For a node $v$ with an agent $r_{i}$ on it, we use $r_{i}$ and $v$ interchangeably to denote the node.
At the end of this phase, each agent has the value of the number of triangles in $G$ i.e., $\mathbf{T}(G)$. In the following theorem, we assemble the list of results we obtained during the 3 phases.
@@
    Each agent $r_{i}$ can calculate the number of triangles in $G$, $\mathbf{T}(G)$, in $O(D\cdot\Delta\cdot\log\lambda)$ rounds,
using $O(\Delta\log n)$ bits of memory per agent.
@@ 5 Applications: Truss Decomposition, Triangle Centrality and Local Clustering Coefficient
In this section, we show some applications of our triangle-counting algorithm via mobile agents in an anonymous graph, namely in solving the truss decomposition problem and computing the triangle centrality metric and local clustering coefficient for each node of the graph. Truss decomposition, and clustering coefficient calculations using mobile agents offer significant advantages in environments where decentralized and adaptive operations are crucial. In scenarios with limited or unreliable communication infrastructure, mobile agents can perform local computations, reducing the need for extensive data transmission. This capability is especially valuable in temporary or ad hoc networks, such as those formed during disaster response, where agents can quickly analyse the network’s structure to identify robust sub-graphs or highly connected areas, ensuring that resources are efficiently allocated. For instance, drones operating in disaster zones can perform truss decomposition to pinpoint areas of strong connectivity, directing rescue efforts where the network of survivors or critical infrastructure is most resilient. In urban environments, mobile agents representing vehicles can compute clustering coefficients to identify highly interconnected regions of a city, enabling more effective traffic management and resource distribution. In the military context, UAVs can employ triangle counting to reveal strategic communication hubs, improving surveillance and ensuring critical areas are secured against potential disruptions.
@@ 5.1 Truss Decomposition
In our algorithm, we use the popular truss decomposition method to first calculate what is called the $trussness$ value for each edge. The $trussness$ values are then used to construct the $k$-truss sub-graph for any value of $k$.
@@
Algorithm for Truss Decomposition: We now propose an algorithm for mobile agents to evaluate $trussness$ for each edge of the graph $G(V,E)$. The algorithm is based on the parallel truss decomposition described in [55]. The $trussness$ values determine a partition (thus an equivalence relation) on $E$, where each class has the edges of $G$ with the same $trussness$ value. Let $\mathbf{t_{k}}$ denote the equivalence class of edges having $trussness=t_{k}$. To find the $k$-truss, $T_{k}$ we construct a sub-graph of $G$ with edges from the equivalence classes $\mathbf{t_{k}},\mathbf{t_{k+1}},\dots,\mathbf{t_{max}}$, where $t_{max}$ is the maximum $trussness$ of any edge in $G$. Mathematically, $T_{k}=\cup_{i=k}^{t_{max}}\mathbf{t_{i}}$. Therefore, by computing the trussness for each edge in $G(V,E)$, we obtain a partition (equivalence classes) of $E$, thereby obtaining the $k$-trusses of $G$ for any $k$ by taking the union of the equivalent classes. Therefore, the $k$-truss decomposition of a graph is equivalent to computing the $trussness$ of each edge in the graph.
$trussness$ values of each edge in $G$ can be computed efficiently using existing serial and parallel algorithms. The serial truss decomposition algorithm first calculates $support$ for each edge and iteratively removes a single edge each time until all the edges of the graph are removed. Each time, the edges are sorted in ascending order and the edge with the lowest $support$ is removed. The removed edge (say $e$) keeps its final $support+2$ as its $trussness$ value. Once $e$ is removed, the $support$ value of the edges that formed the triangle with $e$ is re-evaluated. The remaining edges are once again sorted in order before running another pass. The need to sort the edges in order of $support$ makes the algorithm inherently sequential.
In the parallel version of the algorithm, the sorting condition is relaxed with the use of $h-index$ upgradation. For a set of integers $S$, the $h-index$ of $S$ is defined as the largest number $h$ such that there are at least $h$ elements in $S$ that are equal to and greater than $h$. The $trussness$ of an edge is related to $h-index$, for, the $trussness$ of an edge $e$ can be thought of as the largest $k$ such that it is contained in at least $k$ triangles whose edges have a $trussness$ value of at least $k$. In the algorithm, each edge $e$ in $G$ is initialized to its $support$ as the first approximation to its $trussness$. Now, the $support$ values of all triangles with $e$ as an edge are stored in a set $L$ and its $h-index$ is computed. At each iteration, the $h-index$ of $e$ is updated to the smallest of its current value and the $h-index$ of $L$. The algorithm iteratively updates an edge’s $h-index$ by computing the $h-index$ of all edges that support it, until achieving convergence when no updates would happen. The final $h-index$ of each edge before no further updates happen, provide the $trusness$ value for each edge.
@@
    $r_{i}.edge\_set$ to store the edges incident on a particular node containing $r_{i}$. $r_{i}$ stores only edges $(r_{i},r_{j})$, where $r_{j}>r_{i}$. $edge\_set$ requires $O(\Delta\log n)$ bit of memory, where $\Delta$ is the highest degree of a node in $G$. The ordering of edges ensures that each edge becomes associated with exactly one agent.
@@
    $L$ also associated with each edge in $edge\_set$. $L$ is used to store the values $\min\{h(e^{\prime}),h(e^{\prime\prime})\}$ for each triangle $(e,e^{\prime},e^{\prime\prime})$ containing $e$.
@@
    $scheduled$ to identify the edges in the $edge\_set$ that need further updating in subsequent rounds of the algorithm. Initially, every edge in $edge\_set$ is scheduled ($true$).